OpenStudy (anonymous):

Solve x^2 - 7x = -13

OpenStudy (driftracer305):

x^2-7x+13 = 0... now factorize

OpenStudy (anonymous):

i have to use the quadratic formula thoooo:(

OpenStudy (driftracer305):

lol yup.......

OpenStudy (nirmalnema):

x^2-7x+13=0 \[(-b \pm \sqrt{b^2-4ac}) / 2a\]

OpenStudy (haseeb96):

x^2-7x+13=0 here just use the quadratic formula a=1 ,b=-7 and c=13